Find two possitive numbers that differ by 2 and have a product of 20. I am looking for an answer involving "+/- the square root of 21".
Answer by richwmiller(17219)

a-b=2
a*b=20
a=b+2
(b+2)*b=20
b^2+2b-20=0
complete the square
(b+1)^2=21
your +- sqrt(21) is close.
b+1=+-sqrt(21)
b=-1+-sqrt(21)